I am having a lot of trouble with these three problems. Please help? Solve by elimination/substitution.
12x + 13y = 491
12x + 19y = 629
3x – 4y = -198
x – 3y = -111
x + 2y = 47
x - 3y = 7

Answer by checkley79(3341)

12x + 13y = 491
12x + 19y = 629 SUBTRACT.
-------------------------
-6y=-138
y=-138/-6
y=23 ans.
12x+13*23=491
12x+299=491
12x=491-299
12x=192
y=192/12
y=16 ans.
##########
3x – 4y = -198
x – 3y = -111 multiply by -3 & add.
##############
x + 2y = 47
x - 3y = 7 subtract.
######################